If 3 1/2 pounds of bananas cost $.98, how much would one pound cost?
Answer by christinaaa192(28)

3.5/.98= 1/x--- .98(1)=3.5(x) 3.5(x)=.98 ....35(x)=9.8....350(x)=98.... x=98/350 divide 98/350 to get the answer for x